"Bob is considering buying a home and selling it in one year. At t=0 it will cost him $100,000 to buy the home. At t=1 he will sell it for $150,000. Buying transaction costs (at t=0) are 5% of the purchase price, and selling transaction costs (at t=1) are 8% of the sale price. Note: each time period is a year. Write out the Net Present Value function for Bob s investment for a general annual discount rate i. Plug all of the numbers you can into the function. Sample Answer: NPV(i)= -100 + (5)/(1+i)^1 + 105/(1+i)^2"
Answer #1

At t = 0,

Total Cost = 100,000(1 + 0.05) = $105,000

At t = 1,

Total Amount Received = 150,000(1 - 0.08) = $138,000

NPV = -105,000 + 138,000/(1 + i)